# Minghan Chen  Guikun Chen  Wenguan Wang  Yi Yang

### v1

@@ Appendix C Pseudo Code
Appendix D Failure Cases
In this section, we present failure cases of scene graphs generated by Hydra-SGG on VG150 [68] val. In the first row (${}_{\!}$ 6a-c), Hydra-SGG successfully recognizes the airplane in the image, but the predicted label is ‘plane’. Although ‘plane’ is a plausible label, the ground truth label is ‘airplane’, which leads to all predicted relationships being incorrect since they don’t match the ground truth. However, from a semantic perspective, the model correctly identifies these relationships and even recognizes more fine-grained relationships that are not annotated in the ground truth. For instance, the ground truth only marks the right engine of the airplane, but the model correctly identifies both engines. In addition, the model correctly identifies the logo and windows on the airplane, which are reasonable and correct predictions but are considered incorrect because they are not annotated in the ground truth. In the second row (${}_{\!}$ 6d-f), the ground truth annotation uses a single box to label two women, which is an unreasonable annotation. The model correctly predicts two women and identifies that both women are wearing shirts and shorts, which are annotations absent in the ground truth but are semantically accurate.
Figure 6: Failure cases of Hydra-SGG on VG150 [68] val.
Appendix E Discussion
Limitation Analysis. A significant limitation of the current Hydra-SGG is its inability to predict object and relation categories beyond a predefined closed set. The algorithm can only make predictions for object and relation classes that have been explicitly defined and included in the training dataset. This constraint means that the model lacks the capability to recognize or infer novel object types or relationship categories that were not present during the training phase. Such a limitation restricts the model’s applicability in real-world scenarios, where encountering previously unseen objects or relations is common.

### v2

@@ Appendix B Pseudo Code
Algorithm S1: Hydra-SGG core implementation.
Figure S1: Failure cases of Hydra-SGG on VG150 [75] val.
Appendix C Failure Cases
In this section, we present failure cases of scene graphs generated by Hydra-SGG on VG150 [75] val. In the first row (${}_{\!}$ S1a-c), Hydra-SGG successfully recognizes the airplane in the image, but the predicted label is ‘plane’. Although ‘plane’ is a plausible label, the ground truth label is ‘airplane’, which leads to all predicted relationships being incorrect since they don’t match the ground truth. However, from a semantic perspective, the model correctly identifies these relationships and even recognizes more fine-grained relationships that are not annotated in the ground truth. For instance, the ground truth only marks the right engine of the airplane, but the model correctly identifies both engines. In addition, the model correctly identifies the logo and windows on the airplane, which are reasonable and correct predictions but are considered incorrect because they are not annotated in the ground truth. In the second row (${}_{\!}$ S1d-f), the ground truth annotation uses a single box to label two women, which is an unreasonable annotation. The model correctly predicts two women and identifies that both women are wearing shirts and shorts, which are annotations absent in the ground truth but are semantically accurate.
Appendix D Discussion
Limitation Analysis. A significant limitation of the current Hydra-SGG is its inability to predict object and relation categories beyond a predefined closed set. The algorithm can only make predictions for object and relation classes that have been explicitly defined and included in the training dataset. This constraint means that the model lacks the capability to recognize or infer novel object types or relationship categories that were not present during the training phase. Such a limitation restricts the model’s applicability in real-world scenarios, where encountering previously unseen objects or relations is common. In addition, since the current One-to-Many Relation Assignment rules are manually designed, they may lack flexibility and potentially introduce some noise to the training process.
@@ Appendix D Discussion
As mentioned in the limitations, since the current One-to-Many Relation Assignment rules are manually designed, they may introduce noise during training that affects learning. Therefore, another direction for future work could focus on designing better one-to-many strategies to further improve performance. For instance, we could develop a feature composition module to generate new relation triplets, or we can incorporate a relationship refine module to transfer coarse-grained relations into fine-grained relations.